Operations Report – December 16th

Crew 201 Operations Report 16-Dec-2018

SOL: 1

Name of person filing report: César Augusto Serrano Baza

Non-nominal systems: Generator Autostart

Notes on non-nominal systems: None

Generator (hours run): Around 5:30 pm last night to around 9 am today.

Solar— SOC 81%

Diesel Reading – 95%.

Propane Reading – 23%

Ethanol Free Gasoline – N/A gallons. 4 GALLONS

Water (Axillary tank) – 0 gallons
Water (static tank) – 550 gallons (refilled today).
Axillary to Static tank transfer– no
Gallons transferred: 0
Water in GreenHab – Tank is 100% (refilled today).
Water (loft) – Static to Loft Pump used – yes
Water Meter: 6

Toilet tank emptied: NO

Deimos rover used: no, still not functional
Hours: 0
Beginning charge: –
Ending charge: –
Currently charging: no

Sojourner rover used: ASSIGNED TO DIRECTOR
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: YES
Hours: 63
Beginning charge: 100%
Ending charge: 68%
Currently charging: YES

Opportunity rover used: YES
Hours: 43.9
Beginning charge: 100%
Ending charge: 75%
Currently charging: YES

Curiosity rover used: YES
Hours: 59.6
Beginning charge: 100%
Ending charge: 83%
Currently charging: YES

Notes on rovers: Rovers were used today for the crew training. Deimos has been moved off the campus..

# Hours the ATVs were used today: 0
Notes on ATVs: ATVs were taken out of active service and are now parked away from the station.

HabCar used and why, where? Yes, used for refilling the static tank water and the GreenHab water (1 time each tank) to Hanksville gas station.
CrewCar used and why, where? Last crew took it with them.

General notes and comments: We said goodbye to crew 200.

Summary of internet: Nothing to report. We had some problems getting online with comms. We were sending the information but it seemed that comms was not receiving it on time.

Summary of suits and radios: We had a quick training about how to wear the space suit.

Summary of Hab operations: Nothing to report.

Summary of GreenHab operations: Keep water in the plants. Director planted additional crops.

Summary of ScienceDome operations: We installed a heater in the Dome, working as expected.

Summary of RAMM operations: Director cleaned and organized it today.

Copyright © The Mars Society. All rights reserved. | Main Site